OpenStudy (anonymous):

A car traveling at 48mph overtakes a cyclist who, riding at 12 mph, haf at 3-hr head start. How far from the starting point does the car overtakes the cyclist?

OpenStudy (anonymous):

Let \(t\) be in hours and when the cyclist begins his journey. Position of the car: \(x_\text{car}=v_\text{car}(t-3)\) Position of the cyclist: \(x_\text{cyclist}=v_\text{cyclist}t\) Find out when we have \(x_\text{car}=x_\text{cyclist}\), and then use that to answer the question.

OpenStudy (anonymous):

Do I need to add both speed 48 and 12?

OpenStudy (anonymous):

No. The velocities are independent.

OpenStudy (anonymous):

distance between them=36 miles and relative speed = 48-12=36 so time taken =36/36=1hr

OpenStudy (anonymous):

I don't understand it. Can u please explain to me. I really appreciate it

OpenStudy (anonymous):

so what happen to three hours?

OpenStudy (anonymous):

see the the cyclist started riding at 12 miles per hour before the car and the car started after 3 hours...so in 3 hours distance travelled by cyclist=12*3=36miles

OpenStudy (anonymous):

So u mean in one hour the car overtake the cyclists?

OpenStudy (anonymous):